# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Abstract
Giovanni Zanco has been supported by the FN-F grant “MFGPRE: Mean-field games: regularity, path-dependency, economics”. All the authors are affiliated with the GNAMPA group of INdAM (Istituto Nazionale di Alta Matematica “Francesco Severi”). Part of this work has been carried out first while Daria Ghilli, Fausto Gozzi and Giovanni Zanco were visiting the Institute of Advanced Study at Durham University, and then while Giovanni Zanco was visiting the Department of Applied Physics at Waseda University, Tokyo.
@@ 1 Introduction
However, in this context, all the agents of the economy are typically forward looking and they act following their own objectives, which not necessarily coincide with the ones of the other agents or of the government. It is then interesting to have an understanding if (and how) the decisions of the single agents affect and shape the time-space evolution of the whole economic system.
From the mathematical viewpoint this kind of problems can be studied using the methodological tools of Mean Field Games (MFG from now on), a theory introduced in the last 20 years (starting from the seminal papers of Lasry-Lions [22] and Huang-Caines-Malhamé [20]) which allow to shed some light on the outcome of such complex dynamics.
@@
Hence, the purpose of this paper is to study, from the mathematical point of view a family of MFG arising in modeling how the strategic interactions of small agents can shape the time-space evolution of economic variables.
We assume, as a starting point, that the small agents are homogeneous, move across space maximizing their intertemporal utility from consumption minus a disutility from displacement, and have the possibility to accumulate resources for production (such as human capital). Hence their state variables, at any time $t\geq 0$, are the position $x(t)$ and the human capital $h(t)$. The control variables are the velocity $v(t)$ with which they move in the space, and the allocation of production $s(t)$.
The resulting MFG, even in the simplest case, turns out to be very difficult and quite far from what is covered in the actual literature on the topic (see e.g. the books [10, 9]). The main reasons are the presence of non standard interaction terms, the weaker regularity assumptions on the data (in particular Lipschitz continuity and boundedness of the objective are lost), the state constraints. In particular, concerning the interaction term (which we will denote by $F$) we observe that it takes into account the fact that the spatial aggregation of agents increases the (human) capital $h$, through the so-called spillover effects. In particular, because of such structure, this term in not globally Lipschitz with respect to the distribution. Moreover, the positivity state constraints on the human capital brings the need of treating carefully the HJB and the FP equation proving suitable the estimates near the border where $h=0$.
We study such MFG using the PDE approach, i.e. we consider the associated PDE system of Hamilton-Jacobi-Bellman (HJB) and Fokker-Planck [^6] Some work on Fokker-Planck equations for uncontrolled interacting systems with spatial structure and non-standard interaction term has been carried out in [16] and [29].(FP) equations. Our main result is Theorem 4.4 on the existence of solution of such PDE system. This is a departure point to understand the equilibria of the economic system, which will be the object of future research.
@@ 2 The model and the assumptions
where $(x_{\mu},h_{\nu}),(x_{\nu},h_{n}u)$ are random variables with joint law $\pi$ and $\Gamma(\mu,\nu)$ is the set of all couplings of $\mu$ and $\nu$.
@@ 2.1 The economic model
and, for every probability measure $\mu$ on $\mathbb{R}\times\mathbb{R}_{+}$ with first moment finite (recall that $\mathcal{P}_{1}$ is the set of such measures), we also set
@@
Given the above we eventually define the function $F\colon\mathbb{R}\times\mathcal{P}_{1}\rightarrow\mathbb{R}_{+}$ as
@@ 2.2 The Mean Field Game system
$\begin{cases}-\partial_{t}V(t,x,h)+\rho V(t,x,h)=H_{0}(p)+H_{1}(x,h,\mu(t),DV(t,x,h))+\frac{1}{2}\varepsilon^{2}D^{2}_{xx}V(t,x,h)+\frac{1}{2}\chi^{2}h^{2}D^{2}_{hh}V(t,x,h),\\ \partial_{t}\mu(t)=\frac{1}{2}\varepsilon^{2}D^{2}_{xx}\mu(t)+\frac{1}{2}\chi^{2}D^{2}_{hh}\left(h^{2}\mu(t)\right)-D_{x}\left(D_{p}H_{0}(D_{x}V(t,x,h))\mu(t)\right)-D_{h}\left(D_{p}H_{1}(x,h,\mu(t),D_{h}V(t,x,h))\mu(t)\right),\\ \mu(0)=\mu_{0},\quad V(T,x,h)=0\quad\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}.\end{cases}$ (6)
@@
The Hamiltonians $H_{0}\colon\mathbb{R}\rightarrow\mathbb{R}$ and $H_{1}\colon\mathbb{R}\times\mathbb{R}_{+}\times\mathcal{P}_{1}\times\mathbb{R}$ are given by
@@
$P_{j}(\mu)=\int\left|(y,k)\right|^{j}\mu(\mathrm{d}y,\mathrm{d}k)\quad\text{and}\quad M(\mu)=\int k\mu(\mathrm{d}y,\mathrm{d}k);$
as noted before, we have $M(\mu)>0$ if and only if $\mu(\mathbb{R}\times\{0\})<1$ and $M(\mu)=0$ if and only if $\mu(\mathbb{R}\times\{0\})=1$.
@@
with analogous definition for $\underline{M}(\mu(\cdot))$.
@@ 3.1 The Hamilton-Jacobi-Bellman equation
To study the HJB equation we fix $\mu(\cdot)\in C^{\alpha}([0,T];\mathcal{P}_{1})$ and we consider the problem of choosing $(s(\cdot),v(\cdot))\in\mathcal{K}$ as to maximise $J(t_{0},x_{0},h_{0};v(\cdot),s(\cdot),\mu(\cdot))$ subject to (2) and (3), that is, to
@@
    For every fixed $\mu\in\mathcal{P}_{1}$, $x\mapsto F(x,\mu)$ is bounded; if $M(\mu)>0$ then $x\mapsto F(x,\mu)$ is also bounded away from zero. Moreover $F(x,\mu)$ is bounded uniformly for $\mu$ in any relatively compact set of $\mathcal{P}_{1}$.
@@
    for every fixed $\mu\in\mathcal{P}_{1}$, the function $x\mapsto F(x,\mu)$ is ($C^{2}$ and) Lipschitz, with Lipschitz constant depending on $\mu$. Such constant is uniform on relatively compact subsets of $\mathcal{P}_{1}$.
@@
    $\mu\mapsto F(x,\mu)$ is locally Lipschitz continuous (i.e. in any relatively compact subset of $\mathcal{P}_{1}$), uniformly in $x$.
@@
    $M(\mu)\leq P_{1}(\mu)$
    and relatively compact subsets of $\mathcal{P}_{1}$ are uniformly integrable, the claim follows.
@@
    Let $x$ be fixed and let $\mu,\nu\in\mathcal{P}_{1}$. Then
    $\left|F(x,\mu)\right.$ $-\left.F(x,\nu)\right|=\left|\frac{\int\eta_{1}(|x-y|)k\mu(\mathrm{d}y,\mathrm{d}k)\int\eta_{2}(|x-y|)\nu(\mathrm{d}y,\mathrm{d}k)}{\int\eta_{2}(|x-y|)\mu(\mathrm{d}y,\mathrm{d}k)\int\eta_{2}(|x-y|)\nu(\mathrm{d}y,\mathrm{d}k)}\right.$
    $\phantom{aaaaaaaaaaaaaaaaaaaaa}-\left.\frac{\int\eta_{1}(|x-y|)k\nu(\mathrm{d}y,\mathrm{d}k)\int\eta_{2}(|x-y|)\mu(\mathrm{d}y,\mathrm{d}k)}{\int\eta_{2}(|x-y|)\nu(\mathrm{d}y,\mathrm{d}k)\int\eta_{2}(|x-y|)\mu(\mathrm{d}y,\mathrm{d}k)}\right|$
    $\leq\frac{1}{\theta^{2}}\left|\int\eta_{1}(|x-y|)k\mu(\mathrm{d}y,\mathrm{d}k)\int\eta_{2}(|x-y|)\nu(\mathrm{d}y,\mathrm{d}k)\right.$
    $\phantom{aaaaaaaaaaaaaaaaaaaaa}\left.-\int\eta_{1}(|x-y|)k\nu(\mathrm{d}y,\mathrm{d}k)\int\eta_{2}(|x-y|)\mu(\mathrm{d}y,\mathrm{d}k)\right|$
    $\leq\frac{\Theta}{\theta^{2}}\left(\int k\mu(\mathrm{d}y,\mathrm{d}k)\left|\int\eta_{2}(|x-y|)d(\mu-\nu)\right|+\left|\int\eta_{1}(|x-y|)kd(\mu-\nu)\right|\right)$
    $\leq\frac{\Theta}{\theta^{2}}P_{1}(\mu)L_{\eta_{2}}^{-1}\mathcal{W}_{1}(\mu,\nu)+\frac{\Theta^{2}}{\theta^{2}}\mathcal{W}_{1}(\mu,\nu).$
@@ 3.1.2 Properties of the value function
$\begin{cases}-\partial_{t}V+\rho V=H_{1}(t,x,h,\mu(t),D_{h}V)+\frac{1}{2}h^{2}\chi^{2}D^{2}_{hh}V+\frac{1}{2}\varepsilon^{2}D^{2}_{xx}V+H_{0}(D_{x}V)&\mbox{in }(0,T)\times\mathbb{R}\times\mathbb{R}_{++}\\ V(T,x,h)=0&\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}.\end{cases}$ (10)
@@
We will need some regularity of the first derivatives of $V$ up to the boundary; this is studied later in Proposition 3.3.
@@
    Set $\eta=(1-\gamma)(1-\sigma)$ and
    $M(\mu(\cdot))=\sup_{t\in[0,T]}\int k\mu(\mathrm{d}y,\mathrm{d}k),$
    which is finite by continuity of $\mu(\cdot)$. We have, almost surely,
@@
    $\left|F(x_{\varepsilon,1}(t),\mu(t))^{\gamma(1-\sigma)}-F(x_{\varepsilon,2}(t),\mu(t))^{\gamma(1-\sigma)}\right|\leq\left(\frac{2L_{\eta}\Theta}{\theta^{2}}M(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left|x_{1}-x_{2}\right|^{\gamma(1-\sigma)}.$
@@
    $\left|U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)-U_{\sigma}\left(x_{\varepsilon,2}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)\right|\\ \leq\frac{f\left(h_{\varepsilon,1}\right)^{\eta}}{1-\sigma}\left(\frac{\Theta}{\theta}M(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left(\overline{A}^{1-\sigma}\left(2\frac{L_{\eta}}{\theta}\right)^{\gamma(1-\sigma)}\left|x_{1}-x_{2}\right|^{\gamma(1-\sigma)}+L_{A}^{1-\sigma}\left|x_{1}-x_{2}\right|^{1-\sigma}\right),$
@@
    $\mathbb{E}\left[\int_{t_{0}}^{T}e^{-\rho t}\left|U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)-U_{\sigma}\left(x_{\varepsilon,2}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)\right|\right]\mathrm{d}t\\ \leq\frac{2^{\eta}}{1-\sigma}L_{f}^{\eta}e^{\frac{\eta}{2}C_{2,2}(T-t_{0})}(T-t_{0})h_{1}^{\eta}\left(\frac{\Theta}{\theta}M(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left(\overline{A}^{1-\sigma}\left(2\frac{L_{\eta}}{\theta}\right)^{\gamma(1-\sigma)}\left|x_{1}-x_{2}\right|^{\gamma(1-\sigma)}+L_{A}^{1-\sigma}\left|x_{1}-x_{2}\right|^{1-\sigma}\right).$ (19)
@@
    $\mathbb{E}\left[\int_{t_{0}}^{T}e^{-\rho t}\left|U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)-U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,2}(t),s_{\varepsilon}(t),\mu(t)\right)\right|\mathrm{d}t\right]\\ \leq\frac{1}{1-\sigma}\overline{A}^{1-\sigma}\left(\frac{\Theta}{\theta}M(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}(T-t_{0})}(T-t_{0})\left(\left|h_{1}-h_{2}\right|^{2}+32TC_{1,2}L_{f}^{2}L_{\eta}^{2}h_{2}^{2}e^{C_{2}(T-t_{0})}\left|x_{1}-x_{2}\right|^{2}\right)^{\frac{\eta}{2}}.$ (20)
@@
    $\leq\frac{\overline{A}^{1-\sigma}}{1-\sigma}\left(\frac{\Theta}{\theta}M(\mu)\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}h_{0}^{\eta}e^{\frac{\eta}{2}C_{2}(t_{2}-t_{1})}(t_{2}-t_{1})+\varepsilon.$
@@
    $\leq\delta+\frac{2^{\eta}}{1-\sigma}e^{\frac{\eta}{2}C_{2,2}T}\mathbb{E}\left[\tilde{h}^{2}\right]^{\frac{\eta}{2}}L_{f}^{\eta}T\left(\frac{\Theta}{\theta}M(\mu)\right)^{\gamma(1-\sigma)}$
@@
    $+\frac{\overline{A}^{1-\sigma}}{1-\sigma}\left(\frac{\Theta}{\theta}M(\mu)\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}T}T$
@@
    $\leq\delta+\frac{2^{1+\eta}}{1-\sigma}e^{\eta C_{2,2}T}h_{0}^{\eta}L_{f}^{\eta}T\left(\frac{\Theta}{\theta}M(\mu)\right)^{\gamma(1-\sigma)}$
@@
    $+\frac{\overline{A}^{1-\sigma}}{1-\sigma}\left(\frac{\Theta}{\theta}M(\mu)\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}T}T$
@@
    as the constants appearing in this laast formula can be bounded uniformly for $(x_{0},h_{0})$ in bounded set, continuity of $V$ with respect to $t$, locally uniformly with respect to $(x.h)$ follows from arbitrariness of $\delta$ and $\varepsilon$. The joint continuity in $(t,x,h)$ is then a consequence of Dini’s Theorem.
@@
$\begin{cases}-\partial_{t}V+\rho V=H_{1}(t,x,h,\mu(t),D_{h}V)+\frac{1}{2}\chi^{2}h^{2}D^{2}_{hh}V+\frac{1}{2}\varepsilon^{2}D^{2}_{xx}V+H_{0}(D_{x}V)&\mbox{in }(0,T)\times\mathbb{R}\times\mathbb{R}_{++}\\ V(T,x,h)=\phi(x,h)\quad\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}\\ V(t,x,0)=0\quad\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (25)
@@
$\begin{cases}\partial_{t}v=\frac{1}{2}\varepsilon^{2}D^{2}_{xx}v+\frac{1}{2}\chi^{2}h^{2}D^{2}_{hh}v+\rho v+H_{0}(D_{x}v)+H_{1}(T-t,x,h,\mu(T-t),D_{h}v)&\mbox{ in }(0,T)\times\mathbb{R}\times\mathbb{R}_{++}\\ v(0,x,h)=\phi(x,h)\quad\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}\\ v(t,x,0)=0\quad\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (26)

### v2

@@ Abstract
Giovanni Zanco has been supported by the F-NF grant “MFGPRE: Mean-field games: regularity, path-dependency, economics”.
All the authors are affiliated with the GNAMPA group of INdAM (Istituto Nazionale di Alta Matematica “Francesco Severi”). Part of this work has been carried out first while the authors were visiting the Institute of Advanced Study at Durham University, and then while Giovanni Zanco was visiting the Department of Applied Physics at Waseda University, Tokyo.
@@ 1 Introduction
However, in this context, all the agents of the economy are typically forward looking and they act following their own objectives, which do not necessarily coincide with the ones of the other agents or of the government. It is then interesting to understand if (and how) the decisions of the single agents affect and shape the time-space evolution of the whole economic system.
From the mathematical viewpoint this kind of problems can be studied using the methodological tools of mean field games (MFGs from now on), a theory introduced in the last 20 years (starting from the seminal papers of Lasry-Lions [22] and Huang-Caines-Malhamé [20]) which allow to shed some light on the outcome of such complex dynamics.
@@
Hence, the purpose of this paper is to study, from the mathematical point of view, a family of MFGs arising in modeling how the strategic interactions of small agents can shape the time-space evolution of economic variables.
We assume, as a starting point, that the small agents are homogeneous, move across space maximizing their intertemporal utility from consumption minus a disutility from displacement, and have the possibility to accumulate resources for production (such as human capital). Hence their state variables, at any time $t\geq 0$, are the position $x(t)$ and the human capital $h(t)$. The control variables are the velocity $v(t)$ with which they move in the space, and the allocation of human capital production $s(t)$.
The resulting MFG, even in the simplest case, turns out to be very difficult and quite far from what is covered in the actual literature on the topic (see e.g. the books [10, 9]). The main reasons are the presence of non standard interaction terms, the weaker regularity assumptions on the data (in particular Lipschitz continuity and boundedness of the objective are lost), the state constraints. In particular, concerning the interaction term (which we will denote by $F$), we observe that it takes into account the fact that the spatial aggregation of agents increases the (human) capital $h$, through the so-called spillover effects. In particular, because of such structure, this term in not globally Lipschitz with respect to the distribution. Moreover, the positivity state constraints on the human capital brings the need of treating carefully the HJB and the FP equation proving suitable estimates near the border where $h=0$.
We study such MFG using the PDE approach, i.e. we consider the associated PDE system of Hamilton-Jacobi-Bellman (HJB) and Fokker-Planck [^5] Some work on Fokker-Planck equations for uncontrolled interacting systems with spatial structure and non-standard interaction term has been carried out in [16] and [29].(FP) equations. Our main result is Theorem 4.4 on the existence of solutions of such PDE system. This is a departure point to understand the equilibria of the economic system, which will be the object of future research.
@@ 2 The model and the assumptions
where $(x_{\mu},h_{\mu}),(x_{\nu},h_{\nu})$ are random variables with joint law $\pi$ and $\Gamma(\mu,\nu)$ is the set of all couplings of $\mu$ and $\nu$.
@@ 2.1 The economic model
and, for every probability measure $\mu$ on $\mathbb{R}\times\mathbb{R}_{+}$ with finite second moment (recall that $\mathcal{P}_{2}$ is the set of such measures), we also set
@@
Given the above, we eventually define the function $F\colon\mathbb{R}\times\mathcal{P}_{2}\rightarrow\mathbb{R}_{+}$ as
@@
Remark 2.1.
The interaction term $F$ is actually well-defined on any set of probability measures whose marginal with respect to $h$ has finite expectation; in particular it is well-defined on $\mathcal{P}_{1}$. However since most of our arguments (including some of the properties of $F$ we exploit, see Lemma 3.1) work only in $\mathcal{P}_{2}$, we set our model in $\mathcal{P}_{2}$ from the beginning.
@@ 2.2 The Mean Field Game system
$\begin{cases}-\partial_{t}V(t,x,h)+\rho V(t,x,h)=H_{0}(p)+H_{1}(x,h,\mu(t),DV(t,x,h))+\frac{1}{2}\epsilon^{2}D^{2}_{xx}V(t,x,h)+\frac{1}{2}\chi^{2}h^{2}D^{2}_{hh}V(t,x,h),\\ \partial_{t}\mu(t)=\frac{1}{2}\epsilon^{2}D^{2}_{xx}\mu(t)+\frac{1}{2}\chi^{2}D^{2}_{hh}\left(h^{2}\mu(t)\right)-D_{x}\left(D_{p}H_{0}(D_{x}V(t,x,h))\mu(t)\right)-D_{h}\left(D_{p}H_{1}(x,h,\mu(t),D_{h}V(t,x,h))\mu(t)\right),\\ \mu(0)=\mu_{0},\quad V(T,x,h)=0\quad\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}.\end{cases}$ (6)
@@
The Hamiltonians $H_{0}\colon\mathbb{R}\rightarrow\mathbb{R}$ and $H_{1}\colon\mathbb{R}\times\mathbb{R}_{+}\times\mathcal{P}_{2}\times\mathbb{R}$ are given by
@@
$P_{j}(\mu)=\int\left|(y,k)\right|^{j}\mu(\mathrm{d}y,\mathrm{d}k),\quad M(\mu)=\int k\mu(\mathrm{d}y,\mathrm{d}k)\text{and}\quad M_{2}(\mu)=\int k^{2}\mu(\mathrm{d}y,\mathrm{d}k);$
clearly $M(\mu)\leq M_{2}(\mu)^{\frac{1}{2}}$ and, as noted before, we have $M(\mu)>0$ if and only if $\mu(\mathbb{R}\times\{0\})<1$ and $M(\mu)=0$ if and only if $\mu(\mathbb{R}\times\{0\})=1$.
@@
with analogous definition for $\underline{M}(\mu(\cdot))$. $\overline{M}_{2}(K)$ and $\underline{M}_{2}(K)$ are defined similarly, integrating the function $k^{2}$ in place of the function $k$.
@@ 3.1 The Hamilton-Jacobi-Bellman equation
To study the HJB equation we fix $\mu(\cdot)\in C^{\alpha}([0,T];\mathcal{P}_{1})$, $\alpha\in(0,1)$, and we consider the problem of choosing $(s(\cdot),v(\cdot))\in\mathcal{K}$ as to maximise $J(t_{0},x_{0},h_{0};v(\cdot),s(\cdot),\mu(\cdot))$ subject to (2) and (3), that is, to
@@
    For every fixed $\mu\in\mathcal{P}_{2}$, $x\mapsto F(x,\mu)$ is bounded; if $M(\mu)>0$ then $x\mapsto F(x,\mu)$ is also bounded away from zero. Moreover $F(x,\mu)$ is bounded uniformly for $\mu$ in any relatively compact set of $\mathcal{P}_{2}$.
@@
    for every fixed $\mu\in\mathcal{P}_{2}$, the function $x\mapsto F(x,\mu)$ is ($C^{2}$ and) Lipschitz, with Lipschitz constant depending on $\mu$. Such constant is uniform on relatively compact subsets of $\mathcal{P}_{2}$.
@@
    $\mu\mapsto F(x,\mu)$ is locally Lipschitz continuous (i.e. in any relatively compact subset of $\mathcal{P}_{2}$), uniformly in $x$.
@@
    $M(\mu)\leq P_{1}(\mu)\leq P_{2}(\mu)^{\frac{1}{2}}$
    and relatively compact subsets of $\mathcal{P}_{2}$ are $2$-uniformly integrable, the claim follows.
@@
    Let $x$ be fixed and let $\mu,\nu\in\mathcal{P}_{2}$. Then
[…]
    Using the Rubinstein-Kantorovich characterization of $\mathcal{W}_{1}$ we find
[…]
    $\leq\Theta M(\mu)L_{\eta_{2}}\mathcal{W}_{1}(\mu,\nu)$
    $\leq\Theta M_{2}(\mu)^{\frac{1}{2}}L_{\eta_{2}}\mathcal{W}_{2}(\mu,\nu).$
    Choose now random variables $(y_{\mu},h_{\mu})$ and $(y_{\nu},h_{\nu})$ with law $\mu$ and $\nu$, respectively, and with joint law $\xi\in\Gamma(\mu,\nu)$, where $\xi$ is the coupling of $\mu$ and $\nu$ that attains the infimum in the definition of $\mathcal{W}_{2}$. Then
[…]
    $\leq\Theta\left(\mathbb{E}\left[\left|\eta_{1}(\left|x-y_{\mu}\right|)-\eta_{1}\left(\left|x-y_{\nu}\right|\right)\right|h_{\mu}\right]+\mathbb{E}\left[\eta_{1}\left(\left|x-y_{\mu}\right|\right)\left|h_{\mu}-h_{\nu}\right|\right]\right)$
    $\leq\Theta\left(L_{\eta_{1}}\mathbb{E}\left[\left|y_{\mu}-y_{\nu}\right|^{2}\right]^{\frac{1}{2}}\mathbb{E}\left[h_{\mu}^{2}\right]^{\frac{1}{2}}+\Theta\mathbb{E}\left[\left|h_{\mu}-h_{\nu}\right|^{2}\right]^{\frac{1}{2}}\right)$
    $\leq\Theta\left(L_{\eta_{1}}M_{2}(\mu)^{\frac{1}{2}}+\Theta\right)\left(\int_{\left(\mathbb{R}\times\mathbb{R}_{+}\right)^{2}}\left(\left|y-z\right|^{2}+\left|h-k\right|^{2}\right)\xi\left(\mathrm{d}y,\mathrm{d}h;\mathrm{d}z,\mathrm{d}k\right)\right)^{\frac{1}{2}}$
    $=\Theta\left(L_{\eta_{1}}M_{2}(\mu)^{\frac{1}{2}}+\Theta\right)\mathcal{W}_{2}(\mu,\nu).$
    Thus
    $\left|F(x,\mu)-F(x,\nu)\right|\leq\frac{\Theta}{\theta^{2}}\left(M_{2}(\mu)^{\frac{1}{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}\right)+\Theta\right)\mathcal{W}_{2}(\mu,\nu),$
    and the last assertion follows immediately from $M_{2}(\mu)\leq P_{2}(\mu)$.
@@ 3.1.2 Properties of the value function
$\begin{cases}-\partial_{t}V+\rho V=H_{1}(t,x,h,\mu(t),D_{h}V)+\frac{1}{2}h^{2}\chi^{2}D^{2}_{hh}V+\frac{1}{2}\epsilon^{2}D^{2}_{xx}V+H_{0}(D_{x}V)&\mbox{ in}(0,T)\times\mathbb{R}\times\mathbb{R}_{++}\\ V(T,x,h)=0&\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}.\end{cases}$ (10)
@@
We will also need some regularity of the first derivatives of $V$ up to the boundary; this is studied later in Proposition 3.3.
@@
    Set $\eta=(1-\gamma)(1-\sigma)$; we have, almost surely,
@@
    $\left|F(x_{\varepsilon,1}(t),\mu(t))^{\gamma(1-\sigma)}-F(x_{\varepsilon,2}(t),\mu(t))^{\gamma(1-\sigma)}\right|\leq\left(\frac{2L_{\eta}\Theta}{\theta^{2}}\overline{M}(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left|x_{1}-x_{2}\right|^{\gamma(1-\sigma)}.$
@@
    $\left|U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)-U_{\sigma}\left(x_{\varepsilon,2}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)\right|\\ \leq\frac{f\left(h_{\varepsilon,1}\right)^{\eta}}{1-\sigma}\left(\frac{\Theta}{\theta}\overline{M}(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left(\overline{A}^{1-\sigma}\left(2\frac{L_{\eta}}{\theta}\right)^{\gamma(1-\sigma)}\left|x_{1}-x_{2}\right|^{\gamma(1-\sigma)}+L_{A}^{1-\sigma}\left|x_{1}-x_{2}\right|^{1-\sigma}\right),$
@@
    $\mathbb{E}\left[\int_{t_{0}}^{T}e^{-\rho t}\left|U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)-U_{\sigma}\left(x_{\varepsilon,2}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)\right|\right]\mathrm{d}t\\ \leq\frac{2^{\eta}}{1-\sigma}L_{f}^{\eta}e^{\frac{\eta}{2}C_{2,2}(T-t_{0})}(T-t_{0})h_{1}^{\eta}\left(\frac{\Theta}{\theta}\overline{M}(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left(\overline{A}^{1-\sigma}\left(2\frac{L_{\eta}}{\theta}\right)^{\gamma(1-\sigma)}\left|x_{1}-x_{2}\right|^{\gamma(1-\sigma)}+L_{A}^{1-\sigma}\left|x_{1}-x_{2}\right|^{1-\sigma}\right).$ (19)
@@
    $\mathbb{E}\left[\int_{t_{0}}^{T}e^{-\rho t}\left|U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)-U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,2}(t),s_{\varepsilon}(t),\mu(t)\right)\right|\mathrm{d}t\right]\\ \leq\frac{1}{1-\sigma}\overline{A}^{1-\sigma}\left(\frac{\Theta}{\theta}\overline{M}(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}(T-t_{0})}(T-t_{0})\left(\left|h_{1}-h_{2}\right|^{2}+32TC_{1,2}L_{f}^{2}L_{\eta}^{2}h_{2}^{2}e^{C_{2}(T-t_{0})}\left|x_{1}-x_{2}\right|^{2}\right)^{\frac{\eta}{2}}.$ (20)
@@
    $\leq\frac{\overline{A}^{1-\sigma}}{1-\sigma}\left(\frac{\Theta}{\theta}\overline{M}(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}h_{0}^{\eta}e^{\frac{\eta}{2}C_{2}(t_{2}-t_{1})}(t_{2}-t_{1})+\varepsilon.$
@@
    $\leq\delta+\frac{2^{\eta}}{1-\sigma}e^{\frac{\eta}{2}C_{2,2}T}\mathbb{E}\left[\tilde{h}^{2}\right]^{\frac{\eta}{2}}L_{f}^{\eta}T\left(\frac{\Theta}{\theta}\overline{M}(\mu)\right)^{\gamma(1-\sigma)}$
@@
    $+\frac{\overline{A}^{1-\sigma}}{1-\sigma}\left(\frac{\Theta}{\theta}\overline{M}(\mu)\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}T}T$
@@
    $\leq\delta+\frac{2^{1+\eta}}{1-\sigma}e^{\eta C_{2,2}T}h_{0}^{\eta}L_{f}^{\eta}T\left(\frac{\Theta}{\theta}\overline{M}(\mu)\right)^{\gamma(1-\sigma)}$
@@
    $+\frac{\overline{A}^{1-\sigma}}{1-\sigma}\left(\frac{\Theta}{\theta}\overline{M}(\mu)\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}T}T$
@@
    as the constants appearing in this last formula can be bounded uniformly for $(x_{0},h_{0})$ in bounded sets, continuity of $V$ with respect to $t$, locally uniformly with respect to $(x.h)$, follows from arbitrariness of $\delta$ and $\varepsilon$. The joint continuity in $(t,x,h)$ is then a consequence of Dini’s Theorem.
@@
$\begin{cases}-\partial_{t}V+\rho V=H_{1}(t,x,h,\mu(t),D_{h}V)+\frac{1}{2}\chi^{2}h^{2}D^{2}_{hh}V+\frac{1}{2}\epsilon^{2}D^{2}_{xx}V+H_{0}(D_{x}V)&\mbox{ in}(0,T)\times\mathbb{R}\times\mathbb{R}_{++}\\ V(T,x,h)=\phi(x,h)\quad\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}\\ V(t,x,0)=0\quad\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (25)
@@
$\begin{cases}\partial_{t}v=\frac{1}{2}\epsilon^{2}D^{2}_{xx}v+\frac{1}{2}\chi^{2}h^{2}D^{2}_{hh}v+\rho v+H_{0}(D_{x}v)+H_{1}(T-t,x,h,\mu(T-t),D_{h}v)&\mbox{in }(0,T)\times\mathbb{R}\times\mathbb{R}_{++}\\ v(0,x,h)=\phi(x,h)\quad\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}\\ v(t,x,0)=0\quad\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (26)